Fermi National Laboratory


Accelerator Update - May 30-June 2

Friday May 30
The day shift began with a scheduled power outage at 7 AM so high voltage personnel could switch power from the Kautz Road Substation back to the Master Substation. The power outage lasted about twelve minutes.

At 7:45 AM, much of the lab suffered another power outage. A ComEd high voltage line switch failed. Fermi high voltage technicians had the power switched back to the Kautz Rd. Substation by 8:15 AM.

At 1 PM, water techs discovered a LCW leak in the Tevatron (TeV) at sector C1. They began repairs.

At 4 PM, ComEd reported finding a bad conductor on the power line opposite of the one they were working on and would need another three and half hours to repair the switching tower.

ComEd reported finishing their work on the line switch at 7 PM.

Fermi and ComEd high voltage techs switched site power from the Kautz Rd. Substation to the Master Substation at 9:11 PM. The power outage for this change lasted about ten minutes.

At 11:17 PM, Linac techs noticed a problem with an RF station (KRF1). An expert was called to assist them with repairs.

Operators had both the TeV and the Main Injector (MI) powered by midnight.

Saturday May 31
The midnight shift began with machine experts recovering from the power outage, and with D-zero (D0) in access.

At 1 AM, Operations reported the following machine status: Linac turn on progressing, Booster ready, MI ramping, TeV system restores progressing, Antiproton Source (Pbar) investigating controls and LCW problems.

D0 experimenters came out of access at 2 AM.

Linac techs reported a problem a with RF station (LRF5). An internal LCW leak caused a switch tube failure. The station was back on line by 5:29 AM.

Operations established beam in Linac at 5:30 AM, but getting beam into the Booster was another matter. Operations called in an expert at 6:21 AM.

At 7:13 AM, a Linac RF station (LRF3) developed controls problems. This halted all downstream beam. Operations called a controls expert.

Both the Collider Detector Facility (CDF) and D0 accessed their collision halls a little after 9 AM. They were both out of access by 10:12 AM.

At 11:21 AM, Operation began sending some beam to MiniBooNE.

Operations saw indications of a TeV LCW leak at 12:08 PM. Searches began.

At 12:28 PM, Booster and Linac experts managed to send some beam through to Booster and MiniBooNE.

Operational status at 12:35 PM: experts were tuning up Linac, Booster, Switchyard, and MiniBooNE; MI, and Pbar experts were establishing beam to the Pbar target, and operators continued their search for a TeV LCW leak.

At 1:18 PM, operators found the TeV LCW leak at sector C4. At 1:30 PM, Operations suggested that D0 experimenters check their collision hall for possible water damage. Water techs completed their repairs at 3 PM.

TeV turn on began at 3:21 PM.

At 4:46 PM, the Pbar Lithium Lens tripped off. Operations called in a Pbar expert.

CDF and D0 completed their access at 5:18 PM.

The TeV began circulating beam at 8:19 PM.

At 10 PM, Pbar experts established beam to the Accumulator, which allowed them to resume stacking. Unfortunately, at 11:44 PM, the stack was lost due to an LCW pump problem.

Sunday June 1
The midnight shift began with the Duty Mechanic investigating a pump problem at the Central Utility Building (CUB), with TeV and Switchyard experts tuning, and with MiniBooNE taking beam when available.

At 1 AM, the Duty Mechanic reported finding a shredded coupler for the Pbar LCW pump. They completed their repairs at 2:45 AM.

At 2 AM, Pbar quadrupole shunts were causing problems. Operations called in experts and, later, prepared the area for access.

Pbar experts accessed the Accumulator and had the faulty shunts repaired by 4:34 AM.

Operations established beam to the Accumulator at 5:44 AM. Stacking resumed.

At 10:05 AM, Operations noticed air conditioning problems with H- Source, which wasn't being used as a source. The Duty Mechanic made adjustments to the AC system at 1:18 PM.

Operations put the TeV through a dry squeeze at 1:50 PM. TeV experts resumed tuning at 2:30 PM.

A Linac RF station (LRF5) tripped off at 7:48 PM. It reset okay. It tripped off again at 8:33 PM. It again reset okay, but was not running properly. Operations decided to give it an hour of rest at 8:51 PM.

Operations reestablished Linac beam at 10:12 PM. The MiniBooNE horn had tripped off just prior to this. Operators managed to get the horn back on by 11:50 PM.

At 11:36 PM, Operations put the TeV into shot setup.

Monday June 2
The midnight shift began with the TeV in shot setup, and with MiniBooNE off during the shot due to resting LRF5.

At 1:28 AM, Operations established store 2636 with an initial luminosity of 20.78E30. Operations resumed sending beam to MiniBooNE.

Operations tried to resume stacking, but a LCW problem with the Lithium Lens put a halt to this. Experts were called in and managed to get the pumps on by 3:55 AM. Pbar resumed stacking.

At 5:49 AM, Operators discovered a Booster RF station (BRF7) with a LCW leak. Operators bypassed the station.

A Linac RF station (KRF2) began tripping off at 6:13 AM due to problems with LRF5. Operators adjusted LRF5, which temporarily solved the problems.

Today's Plan
Linacs experts will change out the LRF5 power amplifier tube this morning. During that time, with some things happening after the store is terminated, D0 and CDF will access their collision halls, Recycler experts will access the MI tunnels, and cryo system techs will restart their nitrogen plant and, with permission not yet given, may access the TeV to search for a cryo magnet lead water leak.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/2/2003   email Fermilab